OpenStudy (anonymous):

I need to find the volume of a sqaure pyramid and am give base of 8 and slant height of 13.

OpenStudy (camerondoherty):

Hope this helps: http://bit.do/Pyramid-Volume-Calculator

OpenStudy (camerondoherty):

I can guide you through it though if you would like

OpenStudy (anonymous):

I don't have the height. Is the height 10.25?

OpenStudy (dangerousjesse):

\(\large\frac{hs^2}{3}\)=volume :)

OpenStudy (camerondoherty):

It says slant height of 13...

OpenStudy (anonymous):

but I need height not slant height for volume don't I?

OpenStudy (dangerousjesse):

Length is s and height is h.

OpenStudy (camerondoherty):

Oh...

OpenStudy (anonymous):

what is height?

OpenStudy (camerondoherty):

Slant height is different lol...

OpenStudy (camerondoherty):

Just saw that

OpenStudy (dangerousjesse):

Yes, you're right :) Because you know the sh and base, you can find the height :)

OpenStudy (anonymous):

8^2+x^2=13^2???

OpenStudy (anonymous):

4^2+x^2+13^2

OpenStudy (dangerousjesse):

\[sh=\sqrt{h^2+\frac{s^2}{4}}\] is how you find slant height, so preform the inverse operation to find the height :)

OpenStudy (dangerousjesse):

Legth is s (8)

OpenStudy (anonymous):

I got it height is 12.37

OpenStudy (dangerousjesse):

Yes, that's right :)

OpenStudy (dangerousjesse):

Now \(\large\frac{8\times 12.37^2}{3}\)